Jane plans to retire in 30 years and would like to have $500,000 in her account. If the interest on her account is 3.92% compounded monthly, how much should she put into her account every month?

a. $1,633.33 b. $1,388.89 c. $730.74 d. $653.33


1
Expert's answer
2021-11-08T12:04:29-0500

"A=\frac{P[(1+\frac{r}{n})^{nt}-1]}{\frac{r}{n}}"

A"=" $ 500,000


r"=" "\frac{3.92}{1200}" Compounded monthly


n"=" 12


t"=" 30 years


P"=?"


"\$500,000=\frac{P[(1+\frac{3.92}{1200})^{12×30}-1]}{\frac{3.92}{1200}}"


"\$500,000=684.24P"


"P=\$730.74"


"\therefore" Jane should put "\$" 730.74 into her account every month